A 12 volt battery is connected a light bulb and draws 150mA of current. What is the electrical resistance of a light bulb. How much power does it consume. How much will it cost to operate this bulb for a month,if the cost of electricity is $0.11 per kilo watt

1
Expert's answer
2022-04-28T16:42:36-0400
"R=\frac{V}{I}=\frac{12\:\rm V}{0.150\:\rm A}=80\:\rm \Omega"

"P=VI=12*0.15=1.8\:\rm W"

"\rm cost=1.8/1000*30*24*0.11=0.143\$"